# Enabling Natural Zero-Shot Prompting on Encoder Models via Statement-Tuning

### v2

@@ 4.3 Other Baselines
To assess the feasibility of our approach, we compare Statement-tuned RoBERTa base models with 125 million parameters with a range of competitive multitask fine-tuned encoder-decoder models and decoder-only LLMs spanning a parameter range from 60 million parameters to 7 billion parameters.
We include the following open-source models: Llama-2-7B-chat Touvron et al. (2023), Mistral-7B-Instruct-v0.2 Jiang et al. (2023), QWEN1.5-7B-chat and QWEN1.5-0.5B-chat Bai et al. (2023), Pythia-6.9B and Pythia-2.8B Biderman et al. (2023), Phi-2 Li et al. (2023), FlanT5-Large and FlanT5-Small Chung et al. (2022), and BART-large-mnli Lewis et al. (2020). We use the chat/instruction-tuned version of models to allow for better prompting. We try to select models that have not seen the evaluation data to the best of our knowledge, however, the training data of many of these models is not fully outlined and there can always be the possibility of contamination. We use the standard implementation of all models in the HuggingFace transformers library Wolf et al. (2020).
We train and evaluate all the models on a configuration of 5 AMD EPYC Rome CPU cores and 2 Nvidia Tesla A100 40GB GPUs.
@@ 5.1 Overall Result
Statement-Tuning Enables Effective Zero-Shot Generalization on Encoder Models
Table 1 shows zero-shot performance of statement-tuned RoBERTa. Recall that we explore various statement-tuning sizes, hence here we report the best performance across all training sizes and performance for the 4,000 and 50,000 sample sizes per dataset for the base and large models, respectively. The effect of statement tuning sample size is explored in the later part of this paper.
The result shows that the statement-tuned encoder model can achieve zero-shot generalization across unseen tasks and domains. Our approach achieves respectable performance on unseen balanced COPA, Figurative QA, StoryCloze, Yahoo Topic Answers, and Emotion. Perhaps unsurprisingly, domain generalization in sentiment analysis on Amazon Polarity and paraphrase detection on MRPC is also displayed. We also see that the larger model (RoBERTa-large) achieved much better generalization capability in general.
Comparison Against Larger Zero-Shot Models
Our approach is also competitive against other pretrained open-source encoder-decoder and decoder-only Large Language Models under zero-shot prompting. Despite having significantly fewer parameters than all the models reported (except for FlanT5-small), our approach clearly matches or exceeds many of them on the tasks reported. It is worth noting that our approach with only 125M parameters almost completely outperforms all models under or equal to 6.9B parameters on almost all tasks (except for BCOPA) and is completely dominant on FigQA and StoryCloze, both of which are unrepresented in the training data.
Furthermore, we observe generally stronger performance achieved by the Statement-Tuned RoBERTa-large models with consistently competitive performance that is on par with the 7B parameter LLMs on all the tasks. Notably, the Statement-Tuned RoBERTa-large models greatly outperform all the 7B parameter LLMs (approximately 20 times the number of parameters) on FigQA and StoryCloze, with the best performing RoBERTa-large model scoring an additional 32.6 and 10.7 points over the best performing 7B parameter LLM on the accuracy respectively.
We observe similar results in the 32-shot setting (see Appendix C). These results demonstrate the significant potential of much smaller encoder models being accurate and light alternatives to LLM zero-shot (and few-shot) prompting in natural language understanding.
Figure 3: N-shot accuracy of Statement-Tuned RoBERTa-base models across training datasets of different sizes. The x-axis denotes the number of statements per Statement-Tuning training dataset, with the number of training datasets fixed at 8. Additionally, a baseline comparison with random choice is included for reference.
@@ 5.2 Ablation Studies
Statement Average accuracy
Sample RoB-base RoB-large
1,000 62.2 69.7
2,000 64.7 72.9
3,000 66.2 73.5
4,000 67.3 73.5
5,000 66.9 73.1
10,000 66.2 73.4
20,000 66.7 73.2
40,000 66.5 73.6
50,000 58.0 74.7
@@
Table 2: Average accuracy over all evaluation tasks when trained with different statement sample size.
Statement Finetuning Sample Size
Recall that we only perform statement fine-tuning on a sample of all possible statements from the training dataset. Here, we explore the significance of sample size per-dataset.
As shown in Table 2, a larger sample size does not always mean better overall performance for RoBERTa-base. Specifically, RoBERTa-base’s average performance plateaus around 67.3% after 4k samples and does not increase even with significantly more data. Interestingly, RoBERTa-large benefits and improves significantly from using a larger finetuning size with the best average performance observed when it is finetuned on 50,000 statements per training corpus. We hypothesize that this is due to a larger "capacity" to understand and discriminate between natural language statements which allows RoBERTa-large to benefit from more training data as opposed to RoBERTa-base which has a more limited "capacity" to develop a general semantic understanding of the truthfulness of statements. Due to computational and time constraints, we are not able to exhaust the upper limit of RoBERTa-large’s performance. Despite this, we already achieved remarkable performance with training set sizes of up to 50,000 statements. However, we leave extensive hyperparameter and training set size optimization for future work.
@@
Figure 4: N-shot improvement of Statement-Tuned RoBERTa-base of varying training set sizes across training datasets of different sizes. The y-axis, Delta, is the difference between the accuracy of the Statement-Tuned model and the accuracy achieved by regular-finetuning of RoBERTa-base on the task. A positive Delta indicates improvement over the baseline approach.
@@
While the statement-tuned model shows zero-shot generalization, we can further fine-tune the model with the target downstream task. As seen in Figure 3, we investigate the effect of various statement tuning training sample sizes on the base model’s accuracy across different n-shot continual fine-tuning on the 7 evaluation datasets. We observe a general trend where performance increases with more examples (shots) provided. Amazon Polarity has achieved strong performance even under a zero-shot setting, hence the improvement in a few-shot scenario is limited.
We experimented with re-purposing the model back as a standard classification model. However, performance was not as good on preliminary experimentation so we do not explore this avenue further. Detailed results can be seen in the Appendix D.
N-shot Performance
@@
Additionally, there seems to be a general trend of improved performance by using a larger n for n-shot continual fine-tuning. However, the results seem to indicate a general trend of diminishing returns past using 200-shot finetuning. Nevertheless, it seems that a great deal of the potential performance is achieved with the zero-shot application of the approach, hence further supporting the utility of our approach when task-specific data is scarce. Additionally, as seen in Fig. 5, there appears to be a great degree of correlation between 32-shot and 0-shot model performance, indicating that observed trends in the 0-shot scenario can be informative for the few-shot case.
Effect of Statement Tuning Statement Diversity
As part of our investigation of Statement-Tuning, we would like to explore the effect of template diversity during Statement-Tuning. We hypothesize that randomly applying a larger number of different statement templates per training corpus will allow for improved performance on unseen tasks as it will help the model be more robust to the phrasing of statement templates and prevent it from forming dependencies on superficial cues in some templates.
To recall, each corpus employs several templates (see Appendix A). In this experiment, we limit each corpus to only use the maximum of N different templates, which we call Statements per Category (SPC). We Statement-Tune RoBERTa base models with a fixed training set size of 4,000 statements per corpus with a varying level of SPC.
The result can be seen in Table 3. We observe on average that increasing SPC increases performance to a certain extent before leveling off around 63.7% with an SPC of 4. The result is unsurprising, however, confirms our decision to use many statement templates per task. Detailed exploration of statement design is left to future work.
@@
1 72.4 65.7 59.2 92.1 86.2 33.7 46.8 62.0
2 65.0 61.2 56.9 92.3 79.4 31.9 43.8 58.4
3 69.4 70.7 59.4 91.7 85.4 37.5 48.7 63.5
4 67.6 67.1 58.6 92.5 82.1 38.5 54.8 63.7
5 70.6 71.7 58.9 91.9 79.8 40.3 46.9 63.4
Table 3: The Zero-shot performance of the base model using various degrees of SPC, where a larger SPC value indicates greater statement diversity during training. We report the average as the geometric mean of the task performance to account for the differing accuracy ranges of each task.
@@
Table 4: Comparison of the effect of reducing task diversity in the training of Statement-Tuning models on zero-shot accuracy on unseen datasets. The last common is the average using the geometric mean to account for the different accuracy ranges of the different evaluation sets. The total training set size remains constant at approximately 100,000 statements across all configurations.
Comparison Against Standard Fine-tuning
To observe the improvement over regular fine-tuning of RoBERTa-base, we also include Figure 4, where the y-axis, Delta, represents the improvement over regular fine-tuning for the particular n-shot. For zero-shot, we take random choice as a baseline. Generally, continually fine-tuning our model is better than fine-tuning vanilla RoBERTa under an extremely low N-shot setting. However, in some instances such as BCOPA and (to a certain extent) FigQA, we notice that even for a higher number of few-shot examples we tend to observe a benefit against regular fine-tuning.
Overall, this result demonstrates that the particular strength of our approach tends to be in the few-shot and zero-shot case, whereas, in the case of data availability, it tends to be better to directly fine-tune a RoBERTa-base model on the task. We assume that the superiority of our method in situations with limited target task data can be attributed not only to the improved generalizability gained from multitask statement tuning but also to the data augmentation effect brought by the statement tuning on few-shot target task examples, which allows the model to not only learn the relation between a text and the correct class label but also the wrong ones, thus improving data efficiency.
Effect of Statement Tuning Task Diversity
@@
We explore how important it is to cover a variety of tasks during statement tuning. In this experiment, we group statement tuning datasets across 9 different task categories of Summarization (SU), Sentiment Analysis (SA), Question Answering (QA), Natural Language Inference (NLI), Commonsense Reasoning (CR), Paraphrase Detection (PD), Word Sense Disambiguation (WSD), Intent Classification (IC), and Offensive Language Identification (OLI). Dataset to task categories breakdown can be seen in Appendix E. We then perform statement tuning on RoBERTa-base on various subsets of tasks. To control for the total training size, we dynamically sample the data until we have 100k total statements.
Table 4 shows the 0-shot performance of the statement tuning approach when the training set size is fixed but task types are incrementally removed from the mix. We generally observe that the inclusion of more tasks seems to be overall beneficial; however, it seems to also unsurprisingly correlate with how closely the training set resembles the evaluation set. For example, we see a significant jump in accuracy on Amazon Polarity once we introduce another out-of-domain sentiment analysis data. Likewise, MRPC’s performance is already high since another paraphrase detection dataset is always introduced regardless of the configuration. Nevertheless, for truly unseen tasks, statement task diversity is beneficial, especially for multi-class classification tasks such as Yahoo Answers Topic and Emotion. In average, training task diversity seems to have an overall beneficial effect on generalizability and contributes to the ability of our approach to generalize effectively to unseen tasks.
@@ 6 Conclusion
As part of their emergent abilities, LLMs generalize to many unseen tasks and domains through few-shot and zero-shot prompting, but are prohibitively computationally expensive and difficult to adapt. To address this issue, we investigate Statement-Tuning, a novel technique for few-shot and zero-shot task generalization for encoder models. We find that this approach can match or outperform the few-shot and zero-shot prompting of many much larger decoder-only or encoder-decoder models on many tasks at a fraction of the parameters. Experimentation shows that the approach can be leveraged by training on as few as 16,000 statements. We find Statement-Tuning training task and prompt diversity to be generally helpful. We speculate that the benefits of this approach could extend beyond task generalization and could prove useful for cross-lingual task transfer, and would like to explore this in future work.
@@
While our approach offers advantages in computational efficiency compared to LLMs, the cost scales with the number of possible targets due to the requirement of one forward pass per label. Additionally, task-specific full fine-tuning can still achieve better performance. Furthermore, the effectiveness of our approach in generalizing to unseen tasks relies on the similarity of those tasks to the training data. Tasks with minimal relation to existing labeled datasets might see limited performance compared to tasks with higher similarity. Finally, our reliance on encoder-based models restricts its application to Natural Language Understanding tasks, excluding tasks like translation or summarization.

### v3

@@ 4.3 Other Baselines
To assess the feasibility of our approach, we compare Statement-Tuned RoBERTa base models with 125 million parameters with a range of competitive multitask fine-tuned encoder-decoder models and decoder-only LLMs spanning a parameter range from 60 million parameters to 70 billion parameters. We include the following open-source models: Meta-Llama-3-70B-Instruct AI@Meta (2024), Llama-2-13B-chat, Llama-2-7B-chat Touvron et al. (2023), Mistral-7B-Instruct-v0.2 Jiang et al. (2023), QWEN1.5-7B-chat and QWEN1.5-0.5B-chat Bai et al. (2023), Pythia-6.9B and Pythia-2.8B Biderman et al. (2023), Phi-2 Li et al. (2023), FlanT5-Large and FlanT5-Small Chung et al. (2022), and BART-large-mnli Lewis et al. (2020).
We use the chat/instruction-tuned version of the models to allow for better instruction following. We try to select models that have not seen the evaluation data to the best of our knowledge, however, the training data of many of these models is not fully outlined and there can always be the possibility of contamination Li and Flanigan (2024). Although these models have already been trained on a large number of instruction fine-tuning datasets, to guarantee a fair comparison as much as possible, we additionally instruction-tune a subset of the models using LoRA Hu et al. (2021) on the same training datasets used for Statement-Tuning. [^2] Due to limited computational resources, we are only able to perform this extended training and analysis on a subset of the models ranging from 500M to 13B parameters. Details regarding data formatting, hyper-parameters, and results are reported in Appendix H.
We train and evaluate all the models on a configuration of 5 AMD EPYC Rome CPU cores and at most 4 Nvidia Tesla A100 40GB GPUs (we only use 4 GPUs for inference of the largest LLMs, and 1 GPU for Statement-Tuning).
Figure 3: N-shot accuracy of Statement-Tuned RoBERTa-base models across training datasets of different sizes. The x-axis denotes the number of statements per Statement-Tuning training dataset, with the number of training datasets fixed.
@@ 5.1 Overall Result
Table 1 shows zero-shot performance of statement-tuned RoBERTa and baselines. Recall that we explore various statement-tuning sizes, hence here we report the best performance across all training sizes and performance for the 4,000 and 10,000 sample sizes per dataset for the base and large models, respectively. The effect of statement tuning sample size is explored in Section 5.2.1.
Statement-Tuning Enables Effective Zero-Shot Generalization on Masked Language Models. The result shows that the multitask statement-tuned encoder model can achieve zero-shot generalization across unseen tasks and domains. On BCOPA (unseen task) and Amazon Polarity (unseen domain), our zero-shot statement-tuned models even achieve accuracies on par with models fine-tuned on the full datasets. We also see that the larger model (RoBERTa-large) achieved much better generalization than the base model in general.
Comparison Against Larger Zero-Shot Models. Our approach is also competitive against other pre-trained open-source encoder-decoder and decoder-only Large Language Models under zero-shot prompting. Despite having significantly fewer parameters than all the models reported (except for FlanT5-small), our approach matches or exceeds many of them on the tasks reported. In average, our best Statement-Tuned RoBERTa-large models with only 355M parameters outperform the best performing LLM (Qwen1.5-7B-Chat) by 4.4 and the largest LLM (Meta-Llama-3-70B-Instruct, with approximately 200 times the number of parameters) by 4.8. It is worth noting that our RoBERTa-base models with only 125M parameters almost completely outperforms all models under or equal to 6.9B parameters (except for FlanT5-Large) on all tasks (except for BCOPA). Our models are dominant on FigQA and StoryCloze, both of which are unrepresented in the training data, with the best performing RoBERTa-large model scoring an additional 32.2 and 9.4 points over Llama3-70B-Instruct on the accuracy respectively.
We observe similar results in the 32-shot setting (see Appendix C) and when the LLMs are additionally instruction-tuned on the same data (see Appendix H). These results demonstrate the capabilities of much smaller encoder models as being accurate and light (in terms of parameters; for speed comparison see Appendix G) alternatives to LLM zero-shot (and few-shot) prompting in natural language understanding.
@@ 5.2 Ablation Studies
Statement Sample Average accuracy
RoB-base RoB-large
1,000 63.0 71.1
2,000 62.7 71.0
3,000 63.1 73.3
4,000 65.4 72.9
5,000 64.7 72.2
10,000 64.3 73.4
20,000 64.9 68.6
40,000 64.1 72.0
50,000 58.5 68.2
@@
Table 2: Average accuracy over all evaluation tasks when trained with different statement sample size per dataset.
@@
Figure 4: N-shot improvement of Statement-Tuned RoBERTa-base of varying training set sizes over standard fine-tuning. The y-axis, Delta, is the difference between the accuracy of the Statement-Tuned model and the accuracy achieved by regular fine-tuning of RoBERTa-base on the task. A positive Delta indicates improvement over the baseline approach.
5.2.1 Statement Finetuning Sample Size
Recall that we only perform statement fine-tuning on a sample of all possible statements from the training dataset. Here, we explore the effect of sample size per dataset in the multitask statement fine-tuning on both zero-shot and few-shot performance.
Zero-Shot. As shown in Table 2, with only 1k samples per datasets, we can already reach 96% of the best performance, which is obtained with 4k samples on RoBERTa-base and 10k samples on RoBERTa-large, showing the sample-efficiency of statement-tuning. For RoBERTa-base, introducing more data after 4,000 samples does not further improve the accuracy on downstream tasks and even causes a decrease. RoBERTa-large benefits from a larger fine-tuning data size with the best average performance observed when the statement number per training set increases to 10,000, more than doubling the optimal sample size of RoBERTA-base. We hypothesize that this is due to a larger capacity to understand and discriminate between natural language statements which allows RoBERTa-large to benefit from more training data as opposed to RoBERTa-base, which has a more limited capacity to develop a general semantic understanding of the truthfulness of statements. Nonetheless, we also observe a decrease in RoBERTa-large’s performance after 10k samples. The existence of a point of diminishing returns in both models when it comes to Statement-Tuning training data sizes indicates that too many samples may lead to overfitting to the training tasks, which affects the generalizability. This is consistent with previous observation that excessive instruction tuning can hurt the performance of smaller models (8B and below) on held-out tasks due to more limited model capacity Wei et al. (2022a).
@@ 5.2.1 Statement Finetuning Sample Size
Few-Shot. While the statement-tuned model shows zero-shot generalization, we can further fine-tune the model on the target downstream task. As seen in Figure 3, we investigate the effect of both the multitask statement-tuning sample size and the number of shots from the target tasks on the n-shot performance on the 7 evaluation datasets.
When increasing the multitask statement-tuning sample size, we observe a trend in n-shot performance similar to the general zero-shot performance shown in Table 2. For example, the optimal data size is achieved at around 4k~5k on COPA, Emotion and Yahoo, and there is an apparent drop in accuracy across different shot numbers and tasks when increasing the sample size from 40k to 50k. There turns out to be a high degree of correlation among all n-shot and 0-shot performance (see Figure 6), indicating that observed trends in the 0-shot scenario can be informative for the few-shot cases.
@@
However, the results seem to indicate a general trend of diminishing returns past using 200-shot finetuning. Nevertheless, it seems that a great deal of the potential performance is achieved with the zero-shot application of the approach, hence further supporting the utility of our approach when task-specific data is scarce.
5.2.2 Comparison with Standard Fine-Tuning
To observe the improvement over regular fine-tuning of RoBERTa-base, we also include Figure 4, where the y-axis, Delta, represents the improvement over regular fine-tuning for the particular n-shot. For zero-shot, we take random choice as the baseline. Generally, continually fine-tuning our model is better than fine-tuning vanilla RoBERTa under an extremely low N-shot setting. However, in some instances such as BCOPA and (to a certain extent) FigQA, we notice that even for a higher number of few-shot examples, we tend to observe a benefit against regular fine-tuning.
Our approach is recommended in extreme few-shot and zero-shot scenarios. When more data is available, directly fine-tuning RoBERTa-base is better. Our method’s good performance with limited data can be attributed to improved generalizability from multitask statement tuning and data augmentation effect of statements generated from few-shot examples, which enhances data efficiency.
@@ 5.2.2 Comparison with Standard Fine-Tuning
1 91.0(1.7) 74.8(2.2) 49.8(4.8) 59.1(0.5) 58.4(17.1) 78.0(3.4) 41.2(2.1) 62.6(6.9)
[…]
Table 3: The Zero-shot performance of the base model using various degrees of SPC, where a larger SPC value indicates greater statement diversity during training. We report the average as the geometric mean of the task performance to account for the differing accuracy ranges of each task. Each value is a mean over 5 evaluation runs and we include the standard deviation in the parentheses.
@@
Table 4: Comparison of the effect of reducing task diversity in the training of Statement-Tuning models on zero-shot accuracy on unseen datasets. The last column is the average using the geometric mean to account for the different accuracy ranges of the different evaluation sets. The total training set size remains constant at approximately 100,000 statements across all configurations.
5.2.3 Effect of Statement Diversity
As part of our investigation of Statement-Tuning, we would like to explore the effect of template diversity during Statement-Tuning. We hypothesize that randomly applying a larger number of different statement templates per training corpus will allow for improved performance on unseen tasks, as it will make the model more robust to the phrasing of statement templates and prevent it from relying on superficial cues in certain templates.
In our main experiments, each dataset employs several templates (see Appendix A). In this experiment, we limit each corpus to only use the maximum of N different templates, which we call Statements per Category (SPC). We statement-tune RoBERTa base models with a fixed training set size of 4,000 statements per training corpus with a varying level of SPC.
Table 3 shows that though BCOPA and StoryCloze benefits from a larger SPC, increasing SPC doesn’t always boost average task performance, with the highest being 62.6 at SPC 1. However, average standard deviation drops significantly from 6.9% to 3.6% when SPC increases from 1 to 2, reaching a low of 3.0% at SPC 3. This suggests that more template diversity improves stability and consistency. Therefore, using at least 2 different templates is recommended for better robustness.
5.2.4 Effect of Task Diversity
@@
We examine the importance of task variety in Statement-Tuning. Our Statement-Tuning datasets can be grouped into 9 task categories: Summarization (SU), Sentiment Analysis (SA), Question Answering (QA), Natural Language Inference (NLI), Commonsense Reasoning (CR), Paraphrase Detection (PD), Word Sense Disambiguation (WSD), Intent Classification (IC), and Offensive Language Identification (OLI). See Appendix E for the dataset breakdown. We perform statement tuning on RoBERTa-base with various task subsets, dynamically sampling data to maintain 100k total statements.
Table 4 shows the zero-shot performance of the statement tuning approach with a fixed training set size but varying task types. Average performance increases from 26.0 to 61.2 as the number of tasks is increased from the minimum of 3 to the maximum of 9. Robustness also improves, shown by a decrease in average standard deviation from 4.8% to 2.7%. This demonstrates that increasing training task diversity can boost performance and reduce variance. Unsurprisingly, the inclusion of the Sentiment Analysis task substantially improves the performance on Amazon Polarity from the same task category. Another related task, Emotion, also shows a large increase after adding the SA task. Although the inclusion of SA and WSD hurts the performance on a dissimilar task, Yahoo Answer Topic, the accuracy is recovered after adding the more related Intent Classification task, and reaches the highest 36.8 when training with all tasks. However, the enhancement of downstream tasks does not always come from similar training tasks. More interestingly, adding the QA task leads to a significant jump in the performance of all evaluation tasks. Though Paraphrase Identification is always included in the training, MRPC still benefits from the QA task, reflected by a great improvement of 26.8. Both related and unrelated training tasks can have a positive effect on the downstream tasks, highlighting the value of task diversity in multitask statement-tuning. Sometimes adding an unrelated task causes a performance drop on certain datasets, e.g. StoryCloze after adding IC, but including more tasks alleviates the problem, again confirming the advantage of task diversity.
@@ 6 Conclusion
As part of their emergent abilities, LLMs generalize to many unseen tasks/domains through few-shot and zero-shot prompting, but are prohibitively computationally expensive and difficult to adapt. To address this issue, we investigate Statement-Tuning, a novel technique for few-shot and zero-shot task generalization for encoder models. We find that this approach can match/outperform the few-shot and zero-shot prompting of many much larger decoder-only or encoder-decoder models on many tasks at a fraction of the parameters. Experimentation shows that the approach can be leveraged by training on as few as 16,000 statements. We find training task and statement template diversity to be generally helpful. We speculate that the benefits of this approach could extend beyond task generalization and could prove useful for cross-lingual task transfer, and would like to explore this in future work.
@@
While our approach offers advantages in computational efficiency compared to LLMs, the cost scales with the number of possible targets due to the requirement of one forward pass per label. That being said, it is still possible to apply our method in extreme multi-class classification because we do not have to use all possible statements with all possible labels for training, as our model is trained exactly to generalize to unseen labels by learning the relation of label semantics and the input text. Additionally, task-specific full fine-tuning can still achieve better performance in the presence of more training data. Therefore, we recommend the use of our approach in the low-resource/no-resource scenario. Furthermore, our method can be sensitive to the Statement-Tuning training set size and other hyperparameters, hence some exploration of ideal hyperparameters may be required before employing Statement-Tuned models. In addition, we limit our analysis only to English, it would be interesting to observe whether the technique enables cross-lingual transfer but we leave this to future work. Finally, our reliance on encoder-based models restricts its application to Natural Language Understanding tasks, excluding tasks like translation or abstractive summarization.